### v1

# FRAMES: Boosting LLMs with A Four-Quadrant Multi-Stage Pretraining Strategy

Abstract
Large language models (LLMs) have significantly advanced human language understanding and generation, with pretraining data quality and organization being crucial to their performance. Multi-stage pretraining is a promising approach, but existing methods often lack quantitative criteria for data partitioning and instead rely on intuitive heuristics. In this paper, we propose the novel Four-quadRAnt Multi-stage prEtraining Strategy (FRAMES), guided by the established principle of organizing the pretraining process into four stages to achieve significant loss reductions four times. This principle is grounded in two key findings: first, training on high Perplexity (PPL) data followed by low PPL data, and second, training on low PPL difference (PD) data followed by high PD data, both causing the loss to drop significantly twice and performance enhancements. By partitioning data into four quadrants and strategically organizing them, FRAMES achieves a remarkable 16.8% average improvement over random sampling across MMLU and CMMLU, effectively boosting LLM performance.
FRAMES: Boosting LLMs with
A Four-Quadrant Multi-Stage Pretraining Strategy
Xuemiao Zhang1,4,  Feiyu Duan2,4,  Liangyu Xu4thanks: Equal contribution.,  Yongwei Zhou4,  Sirui Wang3,4thanks: Corresponding author.,  Rongxiang Weng4,  Jingang Wang4,  Xunliang Cai4 1 Peking University 2 Beihang University 3 Tsinghua University 4 Meituan zhangxuemiao@pku.edu.cnduanfeiyu@buaa.edu.cnywzhouphd2018@gmail.com {xuliangyu02, wangsirui, wangjingang02, caixunliang}@meituan.com
1 Introduction
LLMs have significantly advanced human language understanding and generation Touvron et al. (2023); Dubey et al. (2024); Islam and Moushi (2024). The quality and organization of pretraining data are crucial as they directly impact LLM performance. Studies show that LLMs require different data at various pretraining stages Yu et al. (2024), offering key insights for developing more efficient pretraining data organization strategies.
Recent studies partition the pretraining process into multiple stages, allowing models to learn from data with distinct characteristics at each stage, which enhances pretraining efficiencyLiu et al. (2021a); Yıldız et al. (2024); Anonymous (2025). For instance, MSP introduces specific tasks and data at each phase, allowing models to learn language structures and semantics from simple to complex progressively Liu et al. (2021a). However, these methods don’t provide quantitative criteria for partitioning data across stages, often relying on intuitive heuristics. The limitation underscores the need for more systematic approaches to optimize multi-stage pretraining.
(a) Accuracy on MMLU.
(b) Accuracy on CMMLU.
Figure 1: Few-shot performance for 3B models shows FRAMES achieves 15.3% improvement over Random across MMLU, and 18.2% on CMMLU.
In this paper, we propose a novel Four-quadRAnt Multi-stage prEtraining Strategy (FRAMES) to boost LLMs’ performance, which is guided by the principle of organizing the pretraining process into four stages to achieve significant loss drops four times. The principle is derived from two key findings based on quantifiable data metrics.
Specifically, we intuitively use the metric Perplexity (PPL) to partition the data into two parts: high PPL and low PPL. Our first key finding reveals that training on high PPL data first, followed by low PPL data, leads to the loss dropping significantly twice and boosts model performance. Inspired by PDPC Zhang et al. (2025), we also introduce the PPL difference (PD) between strong and weak models as another metric for data partitioning. Our second key finding shows that training on low PD data first, followed by high PD data, similarly causes the pretraining loss to drop significantly twice, thereby boosting model performance.
Based on these two major findings, we establish the principle of organizing the pretraining process into four stages to achieve significant loss reductions four times. Specifically, we partition the data into four quadrants based on PPL and PD: Quadrant 1 ($Q_{1}$) with low PD and low PPL, Quadrant 2 ($Q_{2}$) with high PD and low PPL, Quadrant 3 ($Q_{3}$) with low PD and high PPL, and Quadrant 4 ($Q_{4}$) with high PD and high PPL. Through further analysis and experiments, we determine that the optimal strategy is to reorganize the pretraining data in the $Q3\to Q4\to Q1\to Q2$ sequence. We also implement an inner-to-outer hierarchical gradual transition method to ensure smooth data transitions between stages. Practical evidence shows that FRAMES results in loss dropping significantly four times, enhances the models’ emergent abilities, and boosts their performance.
In summary, our contributions are as follows: (1) We identify two key findings: first, training on high PPL data followed by low PPL data, and second, training on low PD data followed by high PD data, both causing the loss to drop significantly twice and performance enhancements. (2) Guided by the principle of continuous loss reduction, we propose FRAMES, a novel strategy that partitions data into four quadrants and organizes the pretraining process into four stages, resulting in loss dropping significantly four times and boosting model performance. (3) Experiments demonstrate a 16.8% improvement over uniform sampling in average performance across MMLU and CMMLU.
2 Four Quadrant Multi-stage Pretraining Strategy
In this section, we first evaluate the effectiveness of PPL and PD as metrics for data organization in two-stage training (Sections 2.1 and 2.2). Based on these insights, we propose FRAMES with both metrics(Section 2.3), as shown in Figure 2.
2.1 Two-stage pretraining guided by PPL
Figure 2: The overall framework of FRAMES.
Research has explored pre-determining the sequence of pretraining data points based on their characteristics, which helps optimize models to globally optimal solutions Pattnaik et al. (2024); Soviany et al. (2022). A critical aspect of this process is selecting appropriate metrics to organize data effectively, thereby minimizing training loss. Characteristics for text data include length, rare word frequency, and syntactic structure Campos (2021). Although these heuristic methods seem reasonable from a human cognitive perspective, they may not necessarily align with the specific requirements of the model. Thus, data characteristics should be determined using metrics that are perceptible to the model and align with the standards of the target tasks Xu et al. (2020). In pretraining tasks, PPL closely aligns with the self-supervised learning objective (language modeling) and effectively evaluates model-data fit, making it an appropriate metric for organizing data.
Experimental Setting
We extract 500B tokens from a bilingual dataset, with both English and Chinese corpora  [^1] For the details of the dataset, please refer to Section 3.1. We train a 1.3B reference model (RM) on the subset using a random sequence and compute PPL of the subset using the RM. Based on the median PPL of the dataset, we partition the training data into two equal subsets: $\text{A}_{\text{PPL}}^{\text{low}}$ and $\text{A}_{\text{PPL}}^{\text{high}}$. The data within each subset is uniformly distributed. We conduct two-stage training on the 3B model in the sequences $\text{A}_{\text{PPL}}^{\text{low}}\rightarrow\text{A}_{\text{PPL}}^{\text{high}}$ and $\text{A}_{\text{PPL}}^{\text{high}}\rightarrow\text{A}_{\text{PPL}}^{\text{low}}$, and compare the results with those from the random training model. We evaluate model performance using MMLUHendrycks et al. (2020) and CMMLULi et al. (2023).
(a) PPL training loss
(b) PD training loss
Figure 3: Two-stage pretraining losses based on PPL and PD, respectively.
Methods MMLU CMMLU
Random 24.84 25.57
$\text{A}_{\text{PPL}}^{\text{low}}\to\text{A}_{\text{PPL}}^{\text{high}}$ 25.96 25.97
$\text{A}_{\text{PPL}}^{\text{high}}\to\text{A}_{\text{PPL}}^{\text{low}}$ 39.63 42.63
Table 1: Accuracy on MMLU and CMMLU for two-stage pretraining based on PPL with 3B models.
Results
Figure 3(a) illustrates the changes in training loss over time. It is observed that training on high PPL data followed by low PPL data results in significant loss reductions occurring twice, ultimately achieving a lower loss level. Conversely, the reverse setting maintains a higher loss, which remains above that of the Random setting. Table 1 shows the model’s benchmark accuracy for the settings $\text{A}_{\text{PPL}}^{\text{low}}\rightarrow\text{A}_{\text{PPL}}^{\text{high}}$ and $\text{A}_{\text{PPL}}^{\text{high}}\rightarrow\text{A}_{\text{PPL}}^{\text{low}}$. Training on high PPL data followed by low PPL data yields significantly higher performance than random training. In contrast, in the $\text{A}_{\text{PPL}}^{\text{low}}\rightarrow\text{A}_{\text{PPL}}^{\text{high}}$ setting, the model shows only slight improvement. This leads to our first key finding: training first on high PPL data followed by low PPL data can cause the loss to drop significantly twice, ultimately boosting model performance.
2.2 Two-stage pertaining guided by PD
Subsequent analysis (detailed in Section 3.4) reveals that relying solely on PPL as a metric presents issues, as it is not stable across diverse data domains. Inspired by PDPC Zhang et al. (2025), PD between strong and weak models can also reflect the difficulty of samples for the models. Consider two models, the weak model $M_{w}$ and the strong model $M_{s}$, both trained on an identical dataset $D$. For any sample $x$, we calculate PD as follows:
$PD(x)=\frac{PPL_{M_{w}}(x)-PPL_{M_{s}}(x)}{PPL_{M_{w}}(x)}$ (1)
where $PPL_{M_{w}}(x)$ and $PPL_{M_{s}}(x)$ are the perplexity values of the sample $x$ calculated using $M_{w}$ and $M_{s}$, respectively. A low PD indicates similar learning efficiency for both models, while a high PD suggests the sample is more challenging for the weak model.
If checkpoints from earlier and later training stages of the same model are viewed as weak and strong models (with the same parameters but improved performance due to more data in later stages), then data with low PD values pose similar difficulty for both early and late stages, while data with high PD values are more challenging for the model’s early checkpoints. In light of the above analysis, PD emerges as a model-aware difficulty metric that is well-suited for organizing text data.
The distributions of PPL and PD on different domains are analyzed in Appendix B.5, PD exhibits a relatively consistent distribution across varied domains, following a normal distribution with a mean value of 0.3. Compared to PPL, PD offers a better advantage in maintaining data diversity throughout each stage of training.
Experimental Setting
We train a 100M parameter RM and then calculate the PD for each sample using both the 100M and 1.3B RMs, referred to as PD(100M-1.3B). Using the median value of PD across all data, we divide the training dataset into two subsets with equal token counts: $\text{A}_{\text{PD}}^{\text{low}}$ and $\text{A}_{\text{PD}}^{\text{high}}$. We conduct a two-stage training process with $\text{A}_{\text{PD}}^{\text{low}}$ first, followed by $\text{A}_{\text{PD}}^{\text{high}}$, and compare the results to the random setting. We use the evaluation methods described in Section 2.1.
Methods MMLU CMMLU
Random 24.84 25.57
$\text{A}_{\text{PD}}^{\text{low}}\to\text{A}_{\text{PD}}^{\text{high}}$ 26.89 27.20
Table 2: Accuracy on MMLU and CMMLU for two- stage pretraining based on PD with 3B models.
Results
Figure 3(b) shows the training loss changes under the $\text{A}_{\text{PD}}^{\text{low}}\rightarrow\text{A}_{\text{PD}}^{\text{high}}$ setup. The loss initially drops rapidly during the low PD phase, then stabilizes, and decreases further with high PD data, eventually falling below the Random model’s loss. It suggests that the first phase sets a beneficial optimization path for the second, helping avoid local optima. Table 2 shows the model’s accuracy under the $\text{A}_{\text{PD}}^{\text{low}}\rightarrow\text{A}_{\text{PD}}^{\text{high}}$ setting. Notably, it exceeds the Random setting by 2.05% on MMLU and 1.63% on CMMLU, which validates PD as an effective metric. Further, this leads to our second key finding: training first on low PD data followed by high PD data can cause the loss to drop significantly twice, ultimately boosting model performance.
2.3 Four-Quadrant Guided Training Strategy
Inspired by the two key findings about PPL and PD, we establish the principle of organizing pretraining data to achieve significant reductions in training loss. Based on this principle, we introduce a novel Four quadRAnt Multi-stage prEtraining Strategy (FRAMES), which uses PPL and PD to partition data and reorganize the data sequence to make the training loss drop significantly. The process of FRAMES is shown in Algorithm 1.
Specifically, we train two RMs on the target training set $D$, with the strong model $M_{s}$ having more parameters than the weak model $M_{w}$. Both models are trained on data from the same distribution and under identical settings. We compute PPL using $M_{s}$ and PD using both $M_{s}$ and $M_{w}$ for each data point in $D$. We use the two metrics to partition $\mathcal{D}$ into four quadrants, as shown in Figure 4(a). The data from four quadrants can be described as follows: Quadrant $\mathbf{Q_{1}}$ contains data that both $M_{s}$ and $M_{w}$ can fit well; Quadrant $\mathbf{Q_{2}}$ contains data that $M_{s}$ learns well, but $M_{w}$ struggles with. Data from this subset is challenging and requires a higher model capacity to understand them; Quadrant $\mathbf{Q_{3}}$ contains data poorly learned by both $M_{s}$ and $M_{w}$; Quadrant $\mathbf{Q_{4}}$ contains data poorly learned by $M_{s}$ and even worse by $M_{w}$.
(a) Four Quadrants.
(b) Losses of four quadrants.
Figure 4: Pretraining losses of four quadrants and main experiments.
Based on the two key findings about PPL and PD, we should follow these constraints:
$\forall x\in S_{i},\forall y\in S_{i+1},\quad\text{PPL}(x)\geq\text{PPL}(y)$ (2)
$\forall x\in S_{i},\forall y\in S_{i+1},\quad\text{PD}(x)\leq\text{PD}(y)$ (3)
where $S_{i}$ represents stage $i$ of the training process. Equation (2) ensures that the PPL of data in stage $i$ is not less than that in stage $i+1$. Equation (3) ensures that the PD of data in stage $i$ is not greater than that in stage $i+1$.
To efficiently organize data from the four quadrants while adhering to constraints, we use a double-loop approach: first, dividing the training into two main phases based on a specific constraint, then further splitting each main phase into two sub-phases according to another constraint. The approach yields two distinct four-stage training strategies: the first strategy follows the sequence $Q_{3}\rightarrow Q_{4}\rightarrow Q_{1}\rightarrow Q_{2}$, while the second strategy adopts $Q_{3}\rightarrow Q_{1}\rightarrow Q_{4}\rightarrow Q_{2}$. Both strategies break the constraints between the second and third stages. However, experiments in Sections 2.1 and 2.2 show that the loss change from low to high PPL is much greater than from high to low PD. Thus, we prioritize the constraints in Equation (2) over that in Equation (3), making the first four-stage training strategy $\mathbf{Q_{3}\to Q_{4}\to Q_{1}\to Q_{2}}$ the better choice. Furthermore, we implement an inner-to-outer hierarchical gradual transition method to ensure smooth data transitions between stages.
Algorithm 1 FRAMES
1:Input: pretraining data $D$, all training steps $K$, smooth steps $n$
2:Output: trained model $\theta_{K}$
3:Initialize model parameters $\theta_{0}$
4:Train RMs on i.i.d. subset of $D$
5:Calculate PPL and PD for all samples in $D$ using RMs
6:Divide $D$ into 4 quadrants: $Q_{1}$, $Q_{2}$, $Q_{3}$, and $Q_{4}$, based on PPL and PD
7:Inner-layer mixing:
8:for $t=1$ to $\frac{|Q_{3}|+|Q_{4}|}{N}$ do
9:     Sample batch $B_{t}^{34}$ from $Q_{3}$ and $Q_{4}$
10:end for
11:$S_{34}\leftarrow\bigcup_{t=1}^{\frac{|Q_{3}|+|Q_{4}|}{N}}B_{t}^{34}$
12:for $t=1$ to $\frac{|Q_{1}|+|Q_{2}|}{N}$ do
13:     Sample batch $B_{t}^{12}$ from $Q_{1}$ and $Q_{2}$
14:end for
15:$S_{12}\leftarrow\bigcup_{t=1}^{\frac{|Q_{1}|+|Q_{2}|}{N}}B_{t}^{12}$
16:for $k=1$ to $K$ do
17:     Outer-layer mixing:
18:     Sample final batch $B_{k}$ from $S_{34}$ and $S_{12}$
19:     Update model parameters $\theta_{k}$ using batch $B_{k}$
20:end for
Formulation of FRAMES with Stage Transition
Given the data organization sequence $S=[Q_{3},Q_{4},Q_{1},Q_{2}]$, we use the S-shape function from PDPCZhang et al. (2025) to organize the data as a hierarchical two-source data mixing process:
$f(p)=\frac{1}{1+exp{(a(p-0.5))}}$ (4)
where $p\in[0,1]$, represents the completion ratio of the pretraining process. $f(p)$ provides a proportion of data from the first source, and $a$ modulates the steepness of the curve. Unlike PDPC, we merely utilize the S-shape function for smoothing during stage transitions. Therefore, we use a larger $a=35$ instead of $a=10$ as in PDPC, to achieve a steeper function curve, as illustrated in Appendix B.4.
Formally, the inner-layer mixing is expressed as:
$B_{t}^{ij}=\{x\mid x\sim Q_{i}\}_{f(p)\cdot N}\cup\{x\mid x\sim Q_{j}\}_{(1-f(p))\cdot N}$ (5)
where N is the batch size. When $i=3$ and $j=4$, the mixed data from $Q_{3}$ and $Q_{4}$ at step $t$ is denoted as $B_{t}^{34}$; similarly, when $i=1$ and $j=2$, the mixed data from $Q_{1}$ and $Q_{2}$ at step $t$ is denoted as $B_{t}^{12}$. We take the union of the batches from the inner-layer data mixes to form the outer-layer mixed data sources $S_{34}$ and $S_{12}$, respectively:
$S_{34}=\bigcup_{t=1}^{\frac{|Q_{3}|+|Q_{4}|}{N}}B_{t}^{34},\quad S_{12}=\bigcup_{t=1}^{\frac{|Q_{1}|+|Q_{2}|}{N}}B_{t}^{12}$ (6)
Then the outer-layer mixing is then expressed as:
$B_{t}=\{x\mid x\sim S_{34}\}_{f(p)\cdot N}\cup\{x\mid x\sim S_{12}\}_{(1-f(p))\cdot N}$ (7)
where $f(p)$ controls the proportions of each part during the inner and outer-layer mixing processes. Notably, FRAMES only organizes the given data without performing selection, which can serve as the final data preprocessing step before pretraining.
3 Experiments
Metric Order ARC-E ARC-C SciQ HellaSw. PIQA AVG.
- Random 56.5 23.6 85.8 34.2 67.3 53.5
PD $\downarrow$ 54.7$\downarrow$1.8 21.8$\downarrow$1.8 87.1$\uparrow$1.3 33.7$\downarrow$0.5 67.8$\uparrow$0.5 53.0$\downarrow$0.5
PD $\uparrow$ 56.1$\downarrow$0.4 21.3$\downarrow$2.3 86.2$\uparrow$0.4 34.4$\uparrow$0.2 67.6$\downarrow$0.2 53.1$\downarrow$0.4
PPL $\downarrow$ 45.5$\downarrow$11.0 20.6$\downarrow$3.0 71.2$\downarrow$14.6 30.3$\downarrow$3.9 63.7$\downarrow$3.6 46.3$\downarrow$7.2
PPL $\uparrow$ 47.8$\downarrow$8.7 17.9$\downarrow$5.7 72.7$\downarrow$13.1 29.1$\downarrow$5.1 62.4$\downarrow$4.9 46.0$\downarrow$7.5
Qu.Edu $\downarrow$ 57.2$\uparrow$0.7 26.4$\uparrow$2.8 79.4$\downarrow$6.4 31.0$\downarrow$3.2 62.7$\downarrow$4.6 51.3$\downarrow$2.2
Qu.Edu $\uparrow$ 56.8$\uparrow$0.3 26.0$\uparrow$2.4 82.1$\downarrow$3.7 32.0$\downarrow$2.2 64.9$\downarrow$2.4 52.4$\downarrow$1.1
PPL S.R. 56.1$\downarrow$0.4 24.1$\uparrow$0.5 87.8$\uparrow$2.0 33.9$\downarrow$0.3 67.4$\uparrow$0.1 53.9$\uparrow$0.4
PPL S. 56.1$\downarrow$0.4 22.6$\downarrow$1.0 85.5$\downarrow$0.3 34.20.0 67.5$\uparrow$0.2 53.2$\downarrow$0.3
Qu.Edu S.R 54.7$\downarrow$1.8 22.9$\downarrow$0.7 86.2$\uparrow$0.4 33.6$\downarrow$0.6 67.4$\uparrow$0.1 53.0$\downarrow$0.5
Qu.Edu S. 54.5$\downarrow$2.0 22.8$\downarrow$0.8 87.8$\uparrow$2.0 34.0$\uparrow$0.2 67.9$\uparrow$0.6 53.4$\downarrow$0.1
PDPC S. 57.3$\uparrow$0.8 26.6$\uparrow$3.0 87.9$\uparrow$2.1 33.7$\downarrow$0.5 68.0$\uparrow$0.7 54.7$\uparrow$1.2
FRAMES - 62.9$\uparrow$6.4 26.5$\uparrow$2.9 90.5$\uparrow$4.7 38.4$\uparrow$4.2 69.6$\uparrow$2.3 57.6$\uparrow$4.1
Table 3: Downstream tasks results for different settings on 1.3B models. We report accuracy for each task, and the best performances are marked in bold. Abbreviations: AVG. = Average.
Method MMLU CMMLU CEVAL BBH ARC-C PIQA AVG.
Random 27.7 27.5 27.2 27.9 33.7 76.0 36.7
PDPC 35.8$\uparrow$8.1 35.6$\uparrow$8.1 36.1$\uparrow$8.9 25.7$\downarrow$2.2 35.8$\uparrow$2.1 76.3$\uparrow$0.3 40.9$\uparrow$4.2
$Q_{3}\to Q_{1}\to Q_{4}\dashrightarrow Q_{2}$ 25.8$\downarrow$1.9 26.4$\downarrow$1.1 25.5$\downarrow$1.6 28.9$\uparrow$1.0 35.2$\uparrow$1.5 75.4$\downarrow$0.6 36.5$\downarrow$0.2
FRAMES 43.0$\uparrow$15.3 45.7$\uparrow$18.2 44.0$\uparrow$16.8 27.90.0 36.5$\uparrow$2.8 76.9$\uparrow$0.9 45.7$\uparrow$9.0
[…]
(a) Accuracy on CEVAL.
(b) Accuracy on average
Figure 5: Few-shot downstream performance on various benchmarks with respect to pretraining iterations for Random and FRAMES.
Figure 6: Pretraining losses of main results. (a) Comparison of different methods. (b) Details of pretraining loss. The marked numbers indicate the process of the loss decreasing four times.
3.1 Settings
Data Source
For 3B models, our pretraining data is sourced from various domains, including books Gao et al. (2020), blogs Baumgartner et al. (2020), patents Sharma et al. (2019), Common Crawl Penedo et al. (2024), and Wikipedia. It comprises 500B tokens each for Chinese and English, similar to the Matrix dataset Zhang et al. (2024a). For 1.3B models, we randomly select 100B tokens from the SlimPajama dataset Soboleva et al. (2023).
Pretraining Setting
We test on 3B parameter model for the main experiments, with a batch size of 640 and a context window length of 8192. The Adam optimizer is used for training within the Megatron framework. In addition we also validate on a 1.3B model, comparing it with other models based on PPL, PD, and QuRating metrics. More details could be found in Appendix B.1.
Evaluation
We use benchmarks including MMLU Hendrycks et al. (2020), CMMLU Li et al. (2023), BBH Suzgun et al. (2022), and CEVAL Huang et al. (2023) for 3B models, covering tasks that assess multi-domain knowledge and complex reasoning. For 1.3B models, we evaluate the models’ performance on the following benchmarks: ARC-E Clark et al. (2018), ARC-C Clark et al. (2018), SciQ Welbl et al. (2017), HellaSwag Zellers et al. (2019), and PIQA Bisk et al. (2020), which include tasks such as knowledge question answering and commonsense reasoning. We apply in-context learning and select examples based on task characteristics. Standard accuracy serves as the final metric for all tasks.
3.2 Main Results
Table 4 shows our experimental results. FRAMES significantly outperforms the Random baseline, with a 9.0% improvement in average performance. For specific tasks, it achieves notable gains of 15.3% on MMLU and 18.2% on CMMLU. In the 1.3B model, FRAMES also surpasses all single-dimension baselines, with 4.1% improvement in average, and a maximum of 6.4% gain in a single task. These findings validate the effectiveness of FRAMES.
Figure 1 and 5 shows the model’s benchmark performance over training steps. The model remains stable initially, with a significant performance increase after 90K steps. Figure 6 illustrates the training loss, showing four declines that align with benchmark performance improvements. This confirms the strong link between training loss and benchmark performance: when dataset phases are equally divided, lower training loss results in better performance. Additionally, FRAMES effectively enhances the model’s emergent capabilities, allowing it to acquire foundational skills quickly and better understand data patterns in later stages, improving overall performance. Through the loss smoothness analysis in Appendix B.6, we found that the loss curve of FRAMES has the lowest high-frequency energy proportion of only 0.02%, significantly lower than Random and PDPC. This indicates that FRAMES can make the model converge more stably and reduce the impact of gradient fluctuations during training.
We find that the $Q_{3}\to Q_{4}\to Q_{1}\to Q_{2}$ strategy significantly outperforms the $Q_{3}\to Q_{1}\to Q_{4}\to Q_{2}$ strategy, with the latter showing a performance decline in the third stage. This supports our argument that prioritizing the PPL dimension over the PD dimension is beneficial.
3.3 Ablation Study
We investigate the differential effects of data from distinct quadrants and to characterize their intrinsic properties. Figure 4(b) presents the training loss trajectories across different quadrants. The results demonstrate a significant disparity in loss magnitude, with high PPL data ($Q_{3}$ and $Q_{4}$) exhibiting substantially greater loss values compared to their low PPL counterparts ($Q_{1}$ and $Q_{2}$). Furthermore, under equivalent PPL conditions, data with higher PD ($Q_{2}$ and $Q_{4}$) consistently enabled the model to converge to lower final loss values than data with lower PD ($Q_{1}$ and $Q_{3}$). Notably, our analysis reveals that PPL exerts a more substantial influence on training loss than PD.
(a) PPL ablation study.
(b) PD ablation study.
Figure 7: Ablation studies of different combinations.
From the perspective of four quadrants, do the conclusions of Sections 2.1 and 2.3 still hold? We test various quadrant combinations and find that:
Training first on high PPL data followed by low PPL data proves to be a superior strategy
Figure 7(a) illustrates the outcomes for each combination. We see a significant performance boost when transitioning from larger PPL data to smaller ones $Q_{3}\rightarrow Q_{1}$. However, moving in the opposite direction $Q_{1}\rightarrow Q_{3}$ and $Q_{2}\rightarrow Q_{4}$ causes a certain degree of accuracy drop, regardless of whether PD is large or small. Over a longer training period, moving from $Q_{1}\rightarrow Q_{2}$ to $Q_{4}$ also results in a performance decline.
Starting with low PD data followed by high PD data is more effective
As shown in Figure 7(b), when PD increases ($Q_{3}\rightarrow Q_{4}$, $Q_{1}\rightarrow Q_{2}$, and $Q_{3}\rightarrow Q_{2}$), the model performs better on the benchmark. Furthermore, in a longer training phase, transitioning from $Q_{3}\rightarrow Q_{1}\rightarrow Q_{2}$ can further enhance the model’s performance.
More results can be found in Appendix B.7.
3.4 Analysis
Figure 8: Human intuitive quality metric distribution.
Data in Quadrants Without Inherent Human Intuitive Cognition Favorability
We investigate the characteristics of data within the four quadrants by extracting 1,000 samples from each and analyzing their quality using four raters from QuRating. As shown in Figure 8, the quality distributions are similar across all dimensions for samples from each quadrant, which suggests that model-perceived measures like PPL and PD don’t significantly correlate with human cognition-based measures, such as knowledge, quality, and diversity. FRAMES maintains consistent data quality throughout the training process, preventing the model from focusing solely on low-quality data at any stage.
Figure 9: Data distribution across different sources.
PPL and PD Distribution
We extract 230,000 samples from 23 domains within our training dataset, covering areas such as Arxiv, Law, Code, and Math. Using the 1.3B reference model, we calculate the PPL of all samples and present the PPL distributions in Figure 13. Significant differences are observed across various domains, with mean PPL values ranging from 5 to 22 and varying degrees of variance. Data from domains such as Wikipedia typically exhibit lower PPL, whereas data from Reddit show higher PPL. This implies that sorting data by PPL could cause significant shifts in domain representation between initial and later training stages, potentially exposing the model to overly homogeneous data during each training stage. Sachdeva et al. (2024) highlight that data diversity enhances model performance, and an uneven distribution may lead to a decline in performance.
We calculate the PD of the sampled data from the 100M and 1.3B reference models and visualize its distribution across various sources, as shown in Figure 14. We find that the PD distribution is quite similar across different sources. This similarity implies that using PD as a metric for data partitioning allows each training phase to include data from various sources, ensuring data diversity throughout the training process and thereby maintaining pretraining efficiency.
Data Distribution of the Four Quadrants
We extract 1 million samples from each of the four quadrant distributions and visualize them, as shown in Figure 9. We observe that each quadrant contains data from a wide range of sources, indicating that FRAMES effectively ensures data diversity throughout the training process. Notably, there are distinct differences among the quadrants. For instance, $Q_{2}$ contains a large number of domain-specific datasets such as Wikipedia, and StackExchange. This aligns with common cognitive understanding, as these datasets may feature more complex sentence structures and terminology, which large models can learn effectively while small models may struggle to master, thus suitable for learning in the later stages of training. This also explains the excellent performance in knowledge-intensive tasks, as the texts in Wikipedia contain numerous factual knowledge. Additionally, the data distribution in $Q_{3}$ is relatively balanced, which is advantageous for the initial training phase. Presenting data from $Q_{3}$ at the beginning of the training process allows the model to learn from sufficiently diverse data, thereby establishing a strong foundation for subsequent learning stages.
Semantic Properties Analysis
To assess distinct semantic properties within each quadrant’s data, we randomly select 1K samples from each quadrant and devise 10 language-text related traits for GPT-4o evaluation. For simplicity and precision, all traits are formulated as yes-or-no questions, and we calculate the percentage of samples meeting each trait’s criteria. Details of all traits are listed in Appendix D. As shown in Figure 10, the data across all four quadrants exhibit similar semantic properties. In addition, we use T5 Raffel et al. (2023) to obtain dense vectors of the samples and perform dimensionality reduction using t-SNE, as shown in Figure 11. We could observe that the data in the four quadrants do not have obvious distinctions at the semantic level. This indicates no obvious difference in the semantic distribution of data between the initial and later training stages, ensuring a diverse range of data is consistently encountered throughout the training process, and avoiding the collapse of the model into a certain preference, thereby damaging generalization performance.
Figure 10: Semantic feature analysis of different quadrants.
Figure 11: Semantic distribution of different quadrants.
4 Related Works
Multi-stage pretraining has emerged as a pivotal strategy in the development of LLMs, enabling models to better capture and utilize diverse data characteristics by dividing the training process into distinct phasesPavlova (2025); Zhao et al. (2024); Liu et al. (2021b); Tan et al. (2022). Liu et al. propose a multi-stage pretraining method that leverages various granularities of information, significantly boosting model performanceLiu et al. (2021b). Yildiz et al. explore the use of multi-stage prompting to improve translation tasks, demonstrating its effectiveness in enhancing downstream applicationsTan et al. (2022).
In LLM pretraining, data preprocessing is key to ensuring dataset quality. Traditional methods use expert rules to filter low-quality data Raffel et al. (2020); Rae et al. (2021); Laurençon et al. (2022); Computer (2023); Penedo et al. (2024) and remove duplicates Lee et al. (2022); Sorscher et al. (2022); Abbas et al. (2023); Soboleva et al. (2023); Tirumala et al. (2024). These approaches enhance quality but may lack semantic depth. To improve semantic selection, strategies involve using targeted sources or proxy models Wenzek et al. (2020); Xie et al. (2023); Marion et al. (2023); Thakkar et al. (2023); Engstrom et al. (2024); Yu et al. (2024). Classifiers automate selection, like the logistic regression model used by Du et al. (2022), with others employing complex scoring Zhang et al. (2024b); Sachdeva et al. (2024). QuRating Wettig et al. (2024) uses multiple raters for nuanced evaluation.
Current methods focus on selection but overlook aligning data characteristics with learning stages, missing opportunities in data organization and sequencing to boost pretraining effectiveness.
5 Conclusion
In this study, we propose the Four-quadRAnt Multi-stage prEtraining Strategy (FRAMES), a novel approach designed to boost the performance of LLMs by systematically organizing the pretraining process into four stages. Guided by the principle of achieving significant loss reductions four times, FRAMES employs a strategic partitioning of pretraining data into four quadrants based on PPL and PD. The experimental results, showing a 9.03% average improvement over random sampling, underscore the effectiveness of FRAMES in optimizing pretraining data organization.
6 Limitations and Future Works
In this study, we simply attempt to split the data into two parts based on the size of PPL and PD. Future research can be more detailed, such as subdividing a dimension into three parts or more, to explore more training stages. In addition, we plan to expand the scope of work to verify the applicability of the two-stage pretraining method in more language model architectures, such as MambaGu and Dao (2023), and Mixture of ExpertsWang et al. (2024). Although our evaluation criteria are already quite comprehensive, it is still possible to extend to a wider range of evaluations, including more detailed domain-specific or interactive tasks, such as the evaluation of analogy reasoning abilityHu et al. (2023). Despite these potential limitations, we firmly believe that our research provides valuable insights and practical contributions to the academic community.
Appendix A Ethical Considerations
We utilized publicly available training corpora from the internet to train our models, which inevitably included biased or harmful content, raising concerns about the safety of the model-generated content. To mitigate this issue, we prioritized the selection of high-quality datasets and implemented rigorous data-cleansing processes to remove harmful elements. Additionally, considering the computationally intensive nature of LLM training and its potential environmental impact, this paper explores a multi-stage training approach, aiming to enhance resource efficiency and reduce environmental pollution during the training process.
Appendix B Experimental Details
B.1 Experimental settings
We train reference models with 100M and 1.3B parameters on 500B tokens randomly sampled from the collected dataset, utilizing the Llama architecture Touvron et al. (2023). We use the 1.3B model to compute PPL and both the 100M and 1.3B models to calculate PD. For the main experiments, we train a 3B model. We set the batch size to 640 and the context window length to 8192. The initial learning rate is $2\times 10^{-4}$, with a warm-up phase of 375M tokens. We apply cosine learning rate scheduling with a weight decay of 0.1. We use the Adam optimizer to train the model within the Megatron framework Shoeybi et al. (2019). Furthermore, we validate our approach on a 1.3B model. Using the same settings as the 3B model, we also compare it with models trained by sequencing based on the PPL, PD, and QuRating Wettig et al. (2024) metrics.
B.2 Benchmark Accuracy of Two-stage Pretraining Based on PPL
Table 5 shows the model’s benchmark accuracy at various steps for the settings $\text{A}_{\text{PPL}}^{\text{low}}\rightarrow\text{A}_{\text{PPL}}^{\text{high}}$ and $\text{A}_{\text{PPL}}^{\text{high}}\rightarrow\text{A}_{\text{PPL}}^{\text{low}}$. Training on high PPL data followed by low PPL data initially yields lower performance than random training at 30K steps, but accuracy improves significantly in the later phases at 60K and 95K steps. In contrast, in the $\text{A}_{\text{PPL}}^{\text{low}}\rightarrow\text{A}_{\text{PPL}}^{\text{high}}$ setting, the model shows slight improvement initially but eventually aligns with the random sequence results. This leads to our first key finding: training first on high PPL data followed by low PPL data can cause the loss to drop significantly twice, ultimately boosting model performance.
Methods Steps MMLU CMMLU
Random 30K 25.43 25.88
$\text{A}_{\text{PPL}}^{\text{low}}\to\text{A}_{\text{PPL}}^{\text{high}}$ 26.65 26.19
$\text{A}_{\text{PPL}}^{\text{high}}\to\text{A}_{\text{PPL}}^{\text{low}}$ 25.29 25.53
Random 60K 25.94 25.84
$\text{A}_{\text{PPL}}^{\text{low}}\to\text{A}_{\text{PPL}}^{\text{high}}$ 25.53 25.48
$\text{A}_{\text{PPL}}^{\text{high}}\to\text{A}_{\text{PPL}}^{\text{low}}$ 33.35 35.37
Random 95K 24.84 25.57
$\text{A}_{\text{PPL}}^{\text{low}}\to\text{A}_{\text{PPL}}^{\text{high}}$ 25.96 25.97
$\text{A}_{\text{PPL}}^{\text{high}}\to\text{A}_{\text{PPL}}^{\text{low}}$ 39.63 42.63
Table 5: Accuracy on MMLU and CMMLU for two-stage pretraining across different steps based on PPL with 3B models.
B.3 Benchmark Accuracy of Two-stage Pretraining Based on PD
Table 6 shows the model’s accuracy as training steps increase under the $\text{A}_{\text{PD}}^{\text{low}}\rightarrow\text{A}_{\text{PD}}^{\text{high}}$ setting. The model consistently outperforms the Random model at most steps. Notably, at 95K steps, it exceeds the Random setting by 2.05% on MMLU and 1.63% on CMMLU, which validates PD as an effective metric. This leads to our second key finding: training first on low PD data followed by high PD data can cause the loss to drop significantly twice, ultimately boosting model performance.
Methods Steps MMLU CMMLU
Random 30K 25.43 25.88
$\text{A}_{\text{PD}}^{\text{low}}\to\text{A}_{\text{PD}}^{\text{high}}$ 26.65 26.19
Random 60K 25.94 25.84
$\text{A}_{\text{PD}}^{\text{low}}\to\text{A}_{\text{PD}}^{\text{high}}$ 26.23 25.10
Random 95K 24.84 25.57
$\text{A}_{\text{PD}}^{\text{low}}\to\text{A}_{\text{PD}}^{\text{high}}$ 26.89 27.20
Table 6: Accuracy on MMLU and CMMLU for two- stage pretraining based on PD with 3B models.
B.4 S-shape Functions with a=10 and a=35
Figure 12: S-shape functions with a=10 and a=35.
By applying the S-shape function, we gradually decrease the proportion of $Q_{3}$ data while increasing the proportion of $Q_{4}$ data, resulting in the mixed data $S_{34}$. Figure 12 illustrates the two forms of Equation 4 with $a=10$ and $a=35$. In this paper, we adopt the steeper form with $a=35$ for smoothing transitions between different stages. In the early stages of model training, the proportion of data from the first stage is close to 1. As training progresses, this proportion approaches 0. During the mid-training phase, there is a gradual transition between the data from the first and second stages.
Figure 13: PPL distribution across different sources.
Figure 14: PD Distribution across different sources.
(a) $Q_{1}$
(b) $Q_{2}$
(c) $Q_{3}$
(d) $Q_{4}$
Figure 15: Data distribution details. We’ve rearranged the order based on the frequency of different sources appearing in each quadrant.
B.5 Data Distribution
Figure 13 and 14 show the PPL and PD distribution of samples in our training set. The PPL of data from different domains shows significant variations, with mean values ranging from 5 to 22 and varying degrees of variance. Sorting data by PPL may lead to imbalanced domain representation between early and late training stages, exposing the model to overly homogeneous data at each stage and potentially degrading performance. In contrast, the distribution of PD is quite similar across different data sources, ensuring that each training phase includes diverse data, thereby maintaining pretraining efficiency and enhancing model performance.
Figure 15 illustrates the domain distribution of data across the four quadrants, arranged in descending order of proportion. Quadrants 1 and 2 predominantly contain data related to code (sourced from Github and Stack Exchange) and knowledge (sourced from Arxiv and Wikipedia). In contrast, the data in Quadrants 3 and 4 primarily originate from books and the Common Crawl.
B.6 Smoother LOSS Curve from FRAMES
To analyze the training loss curve’s smoothness, we use spectral analysis. We perform a Fast Fourier Transform (FFT) on the loss curve $l[n]$ as $Y[k]=\text{FFT}(l[n])$ to convert it to the frequency domain and compute the Power Spectral Density $PSD[k]=\frac{|Y[k]|^{2}}{N}$ to examine energy distribution across frequencies. By selecting a cutoff frequency $f_{c}$, we divide the spectrum into low and high frequencies and calculate the high-frequency energy proportion $R=\frac{\sum_{f>f_{c}}PSD[k]}{\sum_{k=0}^{N/2}PSD[k]}$. A smaller $R$ indicates a smoother temporal curve with fewer high-frequency components.
Our analysis reveals that the loss curve from FRAMES has the lowest high-frequency energy proportion at 0.02%, significantly lower than that of PDPC and Random. This suggests that FRAMES allows the model to converge more stably, and reduces the impact of gradient fluctuations during training. This improvement is due to our four-stage training strategy, which organizes similar data into consecutive batches, thus stabilizing the model training process.
B.7 Ablation Studies
In Figures 16, 17, 18 and 19, we provide a more detailed presentation of the performance of our ablation experiments across various subsets.
PPL Ordering
During two-stage training, we observed that training in the order from high PPL to low PPL enhances the model’s emergent capabilities. This training strategy shows significant advantages across multiple datasets, such as MMLU and CMMLU. Following the training path from $Q_{3}$ to $Q_{1}$ can significantly improve accuracy on benchmarks. However, if the opposite training order is adopted (i.e., from $Q_{1}$ to $Q_{3}$ or from $Q_{2}$ to $Q_{4}$), the model’s performance is similar to that of a randomly initialized model, and in some cases, even worse. This indicates that the impact of training order on model performance is asymmetric, highlighting the importance of properly arranging the training stages.
In three-stage training, the same trend persists. Specifically, after completing training from $Q_{1}$ to $Q_{2}$ and then switching to $Q_{4}$, following the order of data PPL from small to large, the model’s accuracy on multiple benchmarks is even lower than that of Random.
PD Ordering
From the perspective of two-stage training, the training order from small PD to large PD is beneficial for model training. We found that whether in the same PPL region (e.g., $Q_{3}$ to $Q_{4}$ or $Q_{1}$ to $Q_{2}$) or between different PPL regions (e.g., $Q_{3}$ to $Q_{2}$), the model can ultimately achieve performance superior to that of Random.
This conclusion is also applicable in three-stage training. During the transition from the second stage ($Q_{1}$) to the third stage ($Q_{2}$) (Figure 19), the model’s final accuracy still improved. This further supports the effectiveness of the small PD to large PD order, emphasizing the importance of properly arranging training steps in multi-stage training.
(a) MMLU
(b) CMMLU
(c) CEVAL
(d) BBH
Figure 16: PPL related ablation study (2 quadrants)
(a) MMLU
(b) CMMLU
(c) CEVAL
(d) BBH
Figure 17: PPL related ablation study (3 quadrants)
(a) MMLU
(b) CMMLU
(c) CEVAL
(d) BBH
Figure 18: PD related ablation study (2 quadrants)
(a) MMLU
(b) CMMLU
(c) CEVAL
(d) BBH
Figure 19: PD related ablation study (3 quadrants)
Appendix C Data Cases
Table LABEL:PD_quantiles presents samples extracted from each quadrant. The data in the $Q_{1}$ quadrant primarily consists of samples with low PPL and low PD. These samples exhibit homogeneity and a high degree of duplication. The $Q_{2}$ quadrant mainly contains data with low PPL but high PD, featuring text expressions that conform to standard written language conventions, with samples being relatively longer and more complex. The data in $Q_{3}$ and $Q_{4}$, with high PPL and a mix of low and high PD, often feature creative, informal, or unexpected language that doesn’t follow typical grammar rules or standard phrasing, differing significantly from usual writing or speaking styles.
Table 7: Samples from different quadrants.
Quadrant 1
Sample 1: …The need to practice good self-care doesn’t change in this working environment, but how you accomplish this goal might. Much of Arel’s own self-care regimen needed to be adjusted."I was used to weekly massage and monthly chiropractic care. That was gone," she explains. "I am used to runs and yoga and time to meditate in complete silence. That was gone, too." And although there was an onboard fitness center available, Arel often found that she couldn’t get a workout in during the day and was too tired at the end of the day to be invested in the results …
Sample 2: … I have to ask you, why’d you–wha–wha–why are you peeing right here?Creepy Guy: What?Kumar: I mean… why’d you pee right next to me when you could like, choose that bush, or–?Creepy Guy: Well, this bush looked like I should pee on it. Why are you peeing on it?Kumar: Well, no one was here when I chose this bush.Creepy Guy: Oh, so you get to pee on it and no one else does? Huh?Kumar: No, it’s just–I just–Creepy Guy: This your bush? You have a special bond with this bush?Kumar: No, I just thought that–Creepy Guy: You the king of the forest…
Quadrant 2
Sample 1: … boolean insertventas() String sql "INSERT INTO ventas (id_venta, venFechaventa, venId_cliente, venIdadministrador, venTotalventa) VALUES (NULL, ’" + vent() "’, ’" clasu.getId_usuarios() "’, ’1’, ’" + pnlProductos.totall + "’)"; try con cn.getConnection(); ps = con.prepareStatement(sql); ps.executeUpdate(); return true; catch (SQLException ex) Logger.getLogger(LogicaSql.class…
Sample 2: … In terms of providing shorter stay parking, Bell Street multi storey car park is identified as a long stay car park, and the tariffs are so designed to encourage the use of the facility by all day / half day parkers with more flexible tariffs available at other car parks and the on street spaces around the vicinity allow for parking for up to one hour.I have commented that there is no short term (30 minutes to 2 hour) parking available at the West Bell Street multi-storey car park and the …
Quadrant 3
Sample 1: … Maybe it just sagslike a heavy load. Or does it explode?   by Langston HughesIn 1849, Elizabeth Blackwell became the first woman to graduate from a U.S. medical school in N.Y.In 1864, Rebecca Lee Crumpler became the first black woman to graduate from a U.S. medical school in New England.In 1915, women represented approximately 5% of the physician workforce in the U.S.In 1983, women represented approximately 1/3 of U.S. medical school matriculants.In 1976, Grady doctor and cardiologist Nanette K. Wenger is recognized as one of Time Magazine’s Women of the year, and in 1998 she is named Chief of Cardiology.In 2006, women represented nearly one half of all U.S. medical …
Sample 2: … FILED UNDER SEAL PURSUANT TO PROTECTIVE ORDER rise to a direct infringement claim against it. See, e.g., Akami Techs., Inc. v. Limelight Networks, Inc., 797 F.3d 1020, 1023 (Fed. Cir. 2015) (noting entities are liable for performance they control). The evidence further shows that Badoo Software Limited and Badoo Limited are also intimately involved in Badoo Trading’s creation and ownership of the infringing Bumble application. Ex. A at Recital (c). Badoo Limited’s records indicate that its "principal activity" is "to provide mobile development services to Badoo Software Limited, another entity in the Worldwide Vision group…
Quadrant 4
Sample 1: … I myself should be a castaway.Young’s Literal: but I chastise my body, and bring it into servitude, lest by any means, having preached to others – I myself may become disapproved.As noted earlier, Paul now applies the example from the Greek sports arena directly to himself ("I discipline… I myself") and does so that he might present himself as an example or model for other believers to imitate (cp 1Co 4:16, 11:1, 1Th 1:6, cp Heb 6:12, He 13:7, 3Jn 1:11)…
Sample 2: … (B) of from about 0.1 to about 10.0% w/v of a bioadhesive polymeric stabilizer selected from the group consisting of:(i) polyethylene-polypropylene oxide tri-block co-polymers of the formula;(polyethylene oxide)a -(polypropylene oxide)b -(polyethylene oxide)c wherein PA4 a is 46, 52, 62, 75, 97, 98, 122, or 128; PA4 b is 16, 30, 35, 39, 47, 54, or 67; and PA4 c is 46, 52, 62, 75, 97, 98, 122, or 128;(ii) polyvinyl alcohol,(iii) polyvinyl pyrr…
Table 7: (continued)
Appendix D Prompts for Case Study
The prompt used in Section 3.4 to analyze the linguistic features of data across different qhadrants is as follows.
Prompts for Property Recognition You are a language model training data annotator. Your task is to identify whether the given text possesses the following characteristic: {Property} The text to be annotated is: {text} Please determine whether the given text possesses this characteristic according to the above rules. The output format should be "Because…, my answer is X’." where X must be either "yes" or "no." You should remain objective and refrain from adding any further comments after making your choice. 1. Does the text contain polysemous words? Polysemous words may make understanding more difficult. 2. Does the text use specialized terminology? Specialized terminology may require specific domain knowledge to understand. 3. Does understanding the text require specific cultural background knowledge? Cultural background dependence may increase the complexity of understanding. 4. Does the text require logical reasoning to understand? Logical reasoning adds depth to understanding. 5. Does the text contain elements of humor? Humor may affect the way the text is understood. 6. Does the text explore ethical or moral issues? This may increase the depth of thought. 7. Does the text use complex sentence structures? Complex sentence structures may increase the difficulty of understanding. 8. Does the text contain scientific or technical concepts? These concepts may require specific knowledge to understand. 9. Does the text express obvious emotional tones? Emotional tones may affect the understanding of the text. 10. Does understanding the text require additional background knowledge? Background knowledge requirements may affect the comprehensibility of the text.